### v2

# Text to Band Gap: Pre-trained Language Models as Encoders for Semiconductor Band Gap Prediction

Abstract
We investigate the use of transformer-based language models, RoBERTa, T5, and LLaMA, for predicting the band gaps of semiconductor materials directly from textual representations that encode key material features such as chemical composition, crystal system, space group, number of atoms per unit cell, valence electron count, and other relevant electronic and structural properties. Quantum chemistry simulations such as Density Functional Theory (DFT) provide accurate predictions but are computationally intensive, limiting their feasibility for large-scale materials screening. Shallow machine learning (ML) models offer faster alternatives but typically require extensive data preprocessing to convert non-numerical material features into structured numerical inputs, often at the cost of losing critical descriptive information. In contrast, our approach leverages pretrained language models to process textual data directly, eliminating the need for manual feature engineering. We construct material descriptions in two formats: structured strings that combine key features in a consistent template, and natural language narratives generated using the ChatGPT API. For each model, we append a custom regression head and perform task-specific finetuning on a curated dataset of inorganic compounds. Our results show that finetuned language models, particularly the decoder-only LLaMA-3 architecture, can outperform conventional approaches in prediction accuracy and flexibility, achieving a mean absolute error (MAE) of 0.248 eV, root mean squared error (RMSE) of 0.345 eV, and $R^{2}$ of 0.891, compared to the best shallow ML baseline (XGBoost), which achieved an MAE of 0.318 eV, RMSE of 0.537 eV, and $R^{2}$ of 0.838. Notably, LLaMA-3 achieves competitive accuracy with minimal fine-tuning, suggesting its architecture enables more transferable representations for scientific tasks. This work demonstrates the effectiveness of finetuned language models for scientific property prediction and provides a scalable, language-native framework for materials informatics.
keywords:
American Chemical Society,
[…]
1 Introduction
The band gap of semiconductor materials is a fundamental property that directly impacts their electronic and optical behaviors. This parameter dictates crucial attributes such as conductivity, light absorption, and emission, making it essential for the performance of various electronic, optoelectronic, and photovoltaic devices 1. Therefore, the precise prediction and control of the band gap are vital for optimizing semiconductor applications in these fields2.
Band gaps are determined primarily through experimental methods, with UV-visible absorption spectroscopy and photoluminescence spectroscopy being the most commonly used techniques3. However, these experimental methods can only measure the band gaps of synthesized materials and are not applicable to new materials designed theoretically. On the computational side, Density Functional Theory (DFT) has been the primary tool for studying the electronic structure of materials4, 5. While DFT can accurately simulate electronic properties such as band structures and band gaps, its high computational cost and resource-intensive nature make it less practical for high-throughput material screening, especially for complex systems6.
Machine learning (ML) methods have become powerful tools for addressing the computational challenges of DFT. Shallow ML models, such as Random Forest and Support Vector Regression, are commonly used to predict materials properties like band gaps based on material descriptors 7, 8, 9. These models offer a cost-effective alternative to solving the full quantum mechanical equations, significantly reducing computational overhead. However, these models often struggle with non-numerical features, requiring extensive preprocessing to convert material properties into numerical formats. This reliance on extensive preprocessing and feature engineering not only adds complexity but also risks discarding nuanced or qualitative information, such as symmetry, bonding environments, or textual metadata, that could be valuable for accurate property prediction 10, 11. Recent advances in deep learning, particularly Graph Neural Networks (GNNs) 12, have improved the ability to capture the interconnectivity of atomic structures by representing them as graphs. However, this approach requires the conversion of atomic arrangements into graph representations, which adds another layer of preprocessing. Furthermore, GNNs still face limitations in integrating non-numerical properties, such as compound names, into the training process without additional preprocessing steps. These challenges underscore the need for approaches that can seamlessly handle both numerical and non-numerical features in material property predictions while minimizing complex preprocessing requirements.
Language models offer unique advantages by directly utilizing human-readable text data, eliminating the need for elaborate pre-processing while preserving critical information embedded in material descriptions 13, 14, 15, 16. This streamlines the prediction process. Recent advances in natural language processing, particularly with large language models (LLMs), have introduced transformative possibilities for materials science. This simplifies the prediction process compared to conventional ML approaches, which typically require precise atomic coordinates or extensive preprocessing to generate numerical features. In contrast, LLMs can directly process text-based descriptions. Leveraging this capability, we predict band gap values directly from text-formatted input, bypassing the need for detailed structural data and eliminating complex feature engineering.
Recent advances in natural language processing, particularly with LLMs, have introduced transformative opportunities for materials science. For instance, AlloyBERT demonstrates the potential of transformer-based models to predict material properties from descriptive text 17. Similarly, AMGPT showcases the benefits of using composition-based input strings and finetuned LLMs, enabling accurate and efficient predictions for materials science tasks 18, 19. Additionally, CatBERTa, a RoBERTa-based predictive model, has been developed to predict adsorption energy in catalyst systems 13, 14.
In this study, we explore the use of transformer-based language models, RoBERTa, T5, and LLaMA-3, to predict the band gaps of semiconductor materials directly from textual descriptions. These models enable the direct encoding of structured or natural language representations of materials, such as chemical composition, crystal symmetry, and electronic features, without requiring conventional feature engineering. While pretrained language models possess strong linguistic priors, we emphasize that fine-tuning on domain-specific objectives is essential for adapting them to materials property prediction. We implement task-specific regression heads on top of each model and finetune them on a curated dataset of inorganic compounds. This approach allows models to learn mappings from text-based input to scalar band gap values. These models provide a flexible and generalizable framework for property prediction from textual materials data, extending the application of language models beyond conventional natural language tasks into scientific domains such as materials informatics.
2 Methods
2.1 RoBERTa
RoBERTa (Robustly Optimized BERT Approach) is a transformer-based encoder-only language model that builds upon the BERT architecture by refining pretraining strategies and removing restrictive design choices 20. It is widely recognized for its ability to capture rich contextual relationships between tokens through deep bidirectional self-attention, making it highly effective for a range of natural language processing tasks.
RoBERTa is pretrained using the masked language modeling (MLM) objective, where a subset of tokens is masked and the model learns to predict them based on surrounding context. Unlike BERT, RoBERTa omits the next-sentence prediction task and instead trains on longer, contiguous sequences of text 20. It also introduces dynamic masking, larger batch sizes, and improved sentence packing strategies, which collectively enhance the generality and robustness of the learned representations. Tokenization is performed using a byte-level Byte-Pair Encoding (BPE) tokenizer, which allows the model to effectively handle multilingual text, rare subwords, and non-standard Unicode characters.
In this study, we leverage the RoBERTa-base variant, comprising 12 transformer layers, 12 attention heads per layer, and a hidden dimension of 768. While RoBERTa provides strong pretrained representations, our framework emphasizes the necessity of objective-specific fine-tuning to adapt the model to domain-specific tasks such as band gap prediction. The model’s encoder-only structure offers advantages in processing structured text inputs efficiently, particularly in cases where outputs are scalar values rather than generated sequences.
2.2 T5
T5 (Text-to-Text Transfer Transformer) is a unified encoder-decoder architecture that reframes all NLP tasks—including classification, translation, and regression—as text-to-text problems 21. In this study, we utilize the T5-small variant, which consists of 6 encoder and 6 decoder layers, each with a hidden size of 512 and 8 attention heads. Unlike encoder-only models such as RoBERTa, T5 processes input and output sequences through separate but coordinated transformer stacks, allowing for flexible sequence-to-sequence modeling and richer bidirectional context capture within the encoder.
T5 is pretrained using a span corruption objective, in which contiguous spans of tokens in the input are replaced by unique sentinel tokens, and the model learns to reconstruct the missing content. This span-level objective fosters stronger global reasoning compared to token-level masking and supports robust performance on tasks requiring understanding of longer context windows. Architectural enhancements include the use of relative positional embeddings rather than absolute positions, parameter sharing between input and output embeddings, and simplified layer normalization without additive bias. These design choices collectively reduce model complexity and improve generalization.
Tokenization in T5 is performed using SentencePiece, a subword tokenizer that supports language-agnostic processing and eliminates the need for whitespace-based token segmentation 22, 23. It is particularly useful for scientific or symbolic inputs, as it can encode rare tokens and domain-specific terms with high fidelity. During training, subword regularization introduces stochasticity in tokenization, enhancing robustness by exposing the model to a diverse range of token combinations.
While T5’s encoder-decoder structure is typically leveraged for generative tasks, it is also adaptable for scalar prediction tasks like band gap regression. In our implementation, we repurpose the encoder to process material descriptions and append a custom regression head to the encoder output.
2.3 LLAMA-3
LLaMA-3.2-1B is a lightweight, decoder-only transformer model derived from Meta’s LLaMA-3 family of large language models 24, 25. Designed to provide a strong performance-efficiency trade-off, this version incorporates approximately 1 billion parameters and benefits from structured pruning and knowledge distillation from larger models. The architecture emphasizes both representational depth and inference efficiency, making it suitable for scalable applications beyond conventional text generation tasks.
Unlike encoder-only models like RoBERTa or encoder-decoder frameworks such as T5, LLAMA-3 operates in an autoregressive, decoder-only configuration. Despite its unidirectional nature, LLAMA-3 can learn rich representations through deeply stacked attention layers and efficient architectural enhancements. The model employs several innovations that distinguish it from earlier transformers: it uses RMSNorm for pre-normalization to improve training stability, SwiGLU as a more expressive activation function, and Rotary Positional Embeddings (RoPE) to better generalize to longer sequences 26. In some variants, Grouped Query Attention (GQA) is used to reduce computational overhead while maintaining strong modeling capacity.
LLaMA-3 tokenization is performed using BPE via SentencePiece 22, which enables robust handling of diverse character sets and subword units. Numerals are decomposed into individual digit tokens, and rare or unknown characters are represented using byte-level encodings. These choices ensure flexibility in modeling scientific texts that may include specialized symbols, units, or chemical formulas.
Although LLAMA-3 was originally developed for autoregressive text generation, its modular architecture and robust token-level representation capabilities make it a versatile candidate for regression tasks. The model is particularly well-suited for structured and semi-structured textual input, where consistent formatting allows it to encode domain-relevant relationships. In our context, a custom regression head is appended to the decoder’s final hidden states to adapt the model for scalar band gap prediction. As with other language models in this study, fine-tuning on task-specific objectives is necessary to align the model’s pretrained linguistic knowledge with the specialized needs of materials property prediction.
2.4 Dataset
In this study, we utilized the AFLOW database, a comprehensive open repository for computational materials science that contains extensive information on inorganic crystalline materials and their properties 27, 28. Band gap calculations in AFLOW combine first-principles methods with empirical corrections through an automated workflow. The framework uses VASP to perform DFT calculations with the GGA-PBE functional for standard compounds while applying the GGA+U method for strongly correlated systems containing d- and f-block elements. To address GGA’s tendency to underestimate band gaps, AFLOW employs an empirical correction scheme based on a linear regression model derived from 102 benchmark compounds with known experimental values29, 27, 30. This systematic approach, along with the database’s vast size and rich feature space, makes AFLOW particularly well-suited for ML tasks aimed at material property prediction.
For our specific analysis, we selected a subset of 27,600 materials with band gap values ranging between 0 and 5 eV (inclusive). This range was chosen because it encompasses the most relevant band gap values for semiconductors, which are of particular interest in materials science and electronic applications. The lower bound of 0 eV represents materials with metallic behavior, where there is no electronic band gap. Materials that have band gap higher than 5 eV are insulating materials, which have too large band gaps and do not conduct electricity under normal conditions 31, 3. By focusing on materials within this range of 0-5 eV band gaps, we ensure that our model targets materials with practical applications in electronics and optoelectronics.
The dataset was divided into training, validation, and test sets to ensure reliable evaluation and optimization of the model. Specifically, 10% of the data was reserved for the test set to evaluate the final performance of the model. The remaining 90% was further split into 80% for training and 20% for validation, ensuring sufficient data for model training while retaining a representative validation set. This splitting strategy ensured the distribution of band gap values across all subsets, minimizing sampling bias and enhancing representativeness.
2.5 Text Data Format
To investigate the impact of input data representation on model performance, we employed two formats for encoding material property information as text. The first format consists of structured strings, where material attributes, such as chemical composition, crystal structure, and electronic features, were compiled into a consistent, template-based layout. This format emphasizes uniformity and feature alignment across samples, providing a well-controlled structure for the language models to process.
The second format consists of natural language descriptions generated using OpenAI’s GPT-3.5 Turbo API. The same core features were provided to the API to produce narrative-style descriptions, introducing greater linguistic variability and a more conversational tone. Prompts were configured to ensure descriptions remained within a 512-token limit to maintain compatibility with the tokenizer constraints of RoBERTa, and to accommodate input length limits for T5 and LLaMA-3 as well.
Both formats were applied uniformly across all three models: RoBERTa, T5, and LLaMA-3. The models processed these inputs through their native tokenization pipelines, without additional handcrafted feature engineering. For each model, we appended a custom regression head to enable scalar prediction of band gap values and performed fine-tuning on the downstream task using the corresponding textual inputs.
2.6 Input Features
We carefully selected features that capture both the chemical composition and structural properties of the materials, ensuring a comprehensive understanding of their electronic characteristics, especially the band gap. The selected features include chemical formula, atomic species, valence electron count, crystal symmetry, and magnetic properties, all of which are known to play critical roles in determining the electronic structure and band gap of materials. A complete list of the 23 selected features, categorized by their respective domains, is provided in Table 1.
Table 1: Selected feature list. Each feature is accompanied by a specific description explaining its physical significance and contribution to material characterization.
Feature Description
Compound Chemical formula of the material, representing its chemical composition
Species List of atomic species constituting the material
Composition Proportion of each element in the material
Valence cell (iupac) Total number of valence electrons in the unit cell, calculated according to IUPAC standards
Species pseudopotential Type of atomic pseudopotentials used for calculations
Crystal class Describing the symmetry properties of the crystal
Crystal family Indicating the basic geometric features of the crystal
Crystal system Describing the shape and symmetry of the unit cell
Fractional coordinates Representing the relative positions of atoms in the unit cell
Lattice parameters The edge lengths and angles of the unit cell
Lattice system Describing the basic geometric features of the unit cell
Lattice variation Providing a more detailed description of the lattice
Space group of the structure Describing the symmetry of the crystal
Space group change loose Space group determined under looser conditions for crystal structure relaxation, potentially leading to larger symmetry changes
Space group change tight Space group determined under stricter conditions for structure relaxation, resulting in fewer symmetry changes
Point group orbifold Describing the topological properties of the point group
Point group order Indicating the number of symmetry operations in the point group
Point group structure Describing the geometric features of the point group
Point group type Classifying the symmetry properties of the point group
Magnitude of magnetic moment for each atom Describing the local magnetism of the material
Magnetization intensity of each atom Representing magnetism at the atomic scale
Total magnetization intensity of the entire unit cell Describing the overall magnetism of the material
Density Density of the material
The chemical formula represents the basic building blocks of the material, providing critical information about its stoichiometry and composition. The nature and type of atoms constituting the material greatly influence its electronic properties32, 9, 33. For example, the number of valence electrons of each species is crucial for band gap predictions34. The total number of valence electrons per unit cell helps predict how these electrons will contribute to the band structure, which is directly related to the band gap.
Additionally, we paid particular attention to structural features, including crystal class, family, and system, as well as lattice parameters, which include the dimensions and angles of the unit cell. These factors not only shape the arrangement of atoms and their interactions but also define the symmetry and geometric properties of the crystal, directly influencing the distribution of electronic states within the energy bands 35, 36, 30. The space group and point group information were included to account for the effects of symmetry on electronic states and band splitting. Magnetic properties, such as atom magnetic moments and cell magnetization, were also considered due to their relationship with spin distribution and electronegativity3. We specifically chose properties derived from relaxed structures in which the structural configurations have been optimized to minimize energy and stress, ensuring that the atoms are in their equilibrium positions.
3 Results and Discussion
3.1 Framework
Figure 1: Overview of the proposed band gap prediction framework. a The pipeline starts from the AFLOW dataset, followed by feature selection, dataset preparation, and LLM model training for final band gap prediction. b Two input formats are illustrated. string-based representation using direct feature values and description-based format generated by GPT-3.5 turbo. c Visualization of the finetuning process. The input text undergoes tokenization and embedding through a pretrained LLM, followed by a custom regression head for prediction. d Demonstrates the Transformer encoder and the multi-head attention mechanism with Query (Q), Key (K), and Value (V) operations.
We developed a language model-based framework to predict the band gaps of semiconductor materials using transformer-based language models: RoBERTa, T5, and LLaMA-3. As shown in Figure 1a, we constructed the dataset from the AFLOW database by extracting relevant material features and transforming them into textual formats suitable for language model inputs. Two types of textual representations were used (Figure 1b): a structured string format that followed a fixed template, and a more flexible natural language description generated using the GPT-3.5 Turbo API. These formats enabled us to assess how each model handles both highly regular and semantically rich input styles. Additional examples of both the structured strings and natural language descriptions are provided in the Supporting Information.s
At the core of our approach are transformer-based language models, each differing in architecture and tokenization strategy. While RoBERTa uses an encoder-only structure, T5 adopts an encoder-decoder design, and LLaMA-3 operates with a decoder-only architecture. Each model tokenizes the input text using its native tokenizer, such as byte-level BPE or SentencePiece, and processes the sequence to generate contextual embeddings (Figure 1c). These embeddings are passed through a custom regression head to produce a scalar band gap prediction, as shown in Figure 1d. The design of this framework allows for direct comparison across model types and input formats while minimizing the need for handcrafted features or domain-specific encodings.
3.2 Model Performance
We evaluated the performance of three transformer-based language models, RoBERTa, T5, and LLaMA-3, on the task of predicting semiconductor band gaps from text-based material descriptions. Each model was finetuned with a custom regression head and assessed using two types of input formats: a structured string representation and a natural language description generated via GPT-3.5 Turbo. These models were compared against shallow ML baselines, which also use the same input data format), including support vector regression (SVR), random forest, and XGBoost. The accuracy of the model was quantified using three metrics: The mean absolute error (MAE), the root mean square error (RMSE), and the coefficient of determination ($R^{2}$), as shown in Table 2, with the parity plots visualized in Figure 2.
Table 2: Comparison of model performance across different ML approaches. Each transformer-based model was evaluated using both structured string and description input formats. Best performance per metric is shown in bold.
Model Model Type MAE (eV) RMSE (eV) $R^{2}$
SVR Shallow ML $0.601\pm 0.010$ $0.844\pm 0.008$ $0.600\pm 0.008$
Random Forest Shallow ML $0.385\pm 0.006$ $0.609\pm 0.006$ $0.792\pm 0.005$
XGBoost Shallow ML $0.318\pm 0.005$ $0.537\pm 0.005$ $0.838\pm 0.004$
RoBERTa(string) LLM (Encoder) $0.325\pm 0.006$ $0.447\pm 0.005$ $0.855\pm 0.004$
RoBERTa(description) LLM (Encoder) $0.421\pm 0.007$ $0.590\pm 0.006$ $0.797\pm 0.006$
T5(string) LLM (Encoder-Decoder) $0.301\pm 0.007$ $0.448\pm 0.006$ $0.861\pm 0.005$
T5(description) LLM (Encoder-Decoder) $0.446\pm 0.011$ $0.615\pm 0.008$ $0.759\pm 0.008$
LLaMA-3(string) LLM (Decoder) $\mathbf{0.248\pm 0.006}$ $\mathbf{0.345\pm 0.005}$ $\mathbf{0.891\pm 0.004}$
LLaMA-3(description) LLM (Decoder) $0.335\pm 0.008$ $0.473\pm 0.006$ $0.843\pm 0.054$
Figure 2: Parity plots for band gap predictions across models: a SVR, b XGBoost, c Random Forest, d RoBERTa, e T5, f LLaMA-3
Across all models, inputs in structured string format consistently outperformed descriptive natural language inputs. Among the transformer-based models, LLaMA-3 with structured input achieved the best overall performance, with a mean absolute error of 0.248 eV and R² of 0.891. While RoBERTa and T5 also outperformed the shallow ML models, their performance was lower when using GPT-generated descriptions compared to structured string inputs. However, the ability of all three language models to still learn meaningful correlations from unstructured natural language input underscores a key strength of this approach. Rather than viewing this variability as a limitation, we interpret it as evidence of the models’ capacity to generalize from flexible, human-like descriptions of materials. Notably, even in its smallest size, LLaMA-3 demonstrated strong capacity for property prediction when paired with fine-tuning, indicating that decoder-only architectures can serve as lightweight yet powerful alternatives for scientific regression tasks.
3.3 Layer Freezing Analysis
To assess the extent to which each model’s pretrained representations contribute to band gap prediction, we conducted layer freezing experiments for RoBERTa, T5, and LLAMA-3 using structured string inputs. In each experiment, we progressively froze layers and finetuned only a portion of the model. The goal was to determine how much domain-specific tuning is necessary to achieve high predictive accuracy. Table 3 report the MAE, RMSE, and $R^{2}$ values for each model under various freezing configurations. In all cases, a custom regression head was trained on top of the partially or fully frozen transformer backbone.
Table 3: Comparison of layer freezing strategies across RoBERTa, T5, and LLaMA-3 using structured string inputs. The fully finetuned (non-frozen) results are included for reference. All values are reported as mean $\pm$ standard deviation.
Model Freezing Strategy MAE (eV) RMSE (eV) $R^{2}$
RoBERTa Fully finetuned (no freezing) $0.325\pm 0.006$ $0.447\pm 0.005$ $0.855\pm 0.004$
Freeze first layer $0.328\pm 0.008$ $0.448\pm 0.006$ $0.848\pm 0.005$
Freeze all but final 3 layers $0.388\pm 0.009$ $0.510\pm 0.007$ $0.817\pm 0.006$
Freeze all but final layer $0.509\pm 0.012$ $0.648\pm 0.009$ $0.721\pm 0.009$
Freeze all layers $0.663\pm 0.016$ $0.826\pm 0.011$ $0.563\pm 0.013$
T5 Fully finetuned (no freezing) $0.301\pm 0.007$ $0.448\pm 0.006$ $0.861\pm 0.005$
Freeze first layer $0.350\pm 0.008$ $0.504\pm 0.007$ $0.849\pm 0.006$
Freeze all but final 3 layers $0.367\pm 0.009$ $0.516\pm 0.007$ $0.832\pm 0.006$
Freeze all but final layer $0.598\pm 0.014$ $0.784\pm 0.011$ $0.619\pm 0.011$
Freeze all layers $0.792\pm 0.019$ $0.981\pm 0.013$ $0.420\pm 0.014$
LLaMA-3 Fully finetuned (no freezing) $\mathbf{0.248\pm 0.006}$ $\mathbf{0.345\pm 0.005}$ $\mathbf{0.891\pm 0.004}$
Freeze first layer $0.279\pm 0.007$ $0.426\pm 0.006$ $0.878\pm 0.004$
Freeze all but final 3 layers $0.318\pm 0.008$ $0.474\pm 0.006$ $0.851\pm 0.005$
Freeze all but final layer $0.424\pm 0.010$ $0.576\pm 0.008$ $0.793\pm 0.007$
Freeze all layers $0.716\pm 0.017$ $0.893\pm 0.012$ $0.518\pm 0.013$
Throughout this study, we refer to the original language models, trained on general text corpora prior to any materials-specific adaptation, as “pretrained” models. After supervised training on the band gap prediction task, we refer to the models as “finetuned.”
In all three models, predictive performance improved progressively as more transformer layers were unfrozen, highlighting the importance of task-specific fine-tuning in scientific regression applications. This trend highlights a key insight: while LLMs capture broadly useful representations during pretraining, these representations alone are insufficient to achieve optimal accuracy in domain-specific prediction tasks without additional adaptation. In the case of RoBERTa and T5, the most notable performance gains occurred when the final three layers of the model were unfrozen. This suggests that the highest-level contextual embeddings learned during pretraining are not directly aligned with the band gap target property but can be effectively adapted through partial retraining.
LLaMA-3, on the other hand, exhibited relatively strong performance even when the majority of its layers were frozen. For example, when only the final three layers were unfrozen during fine-tuning, LLaMA-3 achieved $R^{2}$ values comparable to those of fully fine-tuned RoBERTa and T5 models. This observation points to its efficient architecture and inductive biases, which may make its learned representations more transferable to new scientific tasks. In particular, the decoder-only architecture of LLAMA-3, combined with its use of Rotary Positional Embeddings and SwiGLU activation functions, appears to confer greater flexibility in adapting to structured input formats such as the material strings used in this study. Furthermore, the ability of LLAMA-3 to achieve high $R^{2}$ values with only the final layer trainable supports its utility in resource-constrained scenarios, such as edge inference or few-shot scientific learning contexts.
Collectively, these results reinforce the conclusion that while pretrained LLMs provide a valuable foundation, their effective use in materials science and other physical domains requires carefully calibrated fine-tuning strategies. Strategic layer freezing not only reduces computational burden during training but also offers insight into which parts of the model architecture contribute most directly to scientific prediction tasks. This understanding is critical for efficiently deploying transformer-based models in practical, data-limited, or performance-critical settings.
3.4 Embedding Analysis
We conducted embedding space analysis using t-SNE visualizations to investigate how each model organizes material representations before and after fine-tuning. As shown in Figure 3, we extracted the first-token embeddings from the pretrained and finetuned versions of RoBERTa, T5, and LLaMA-3 using structured string inputs, and colored the points by crystal system. In the “pretrained” condition, which corresponds to the original language model weights before finetuning on the band gap prediction task, the embeddings reflect structural signals learned from general language corpora.
Figure 3: t-sne visualizations of embedding analysis with crystal systems. a pretrained RoBERTa, b finetuned RoBERTa, c pretrained T5, d finetuned T5, e pretrained LLAMA-3, f finetuned LLaMa-3
Figure 4: t-SNE visualizations of embedding analysis for bandgap. a pretrained RoBERTa, b finetuned RoBERTa, c pretrained T5, d finetuned T5, e pretrained LLaMa-3, f finetuned LLaMa-3
In the pretrained state, the embeddings of all three models exhibit meaningful clustering based on the crystal system. This structure-aware behavior is likely driven by explicit features in the input, such as crystal class, space group, and lattice parameters, information that correlates with symmetry. However, while these features can separate materials by structural family, they do not offer sufficient signal to accurately predict electronic properties like band gap. Without finetuning, the models lack the task-specific supervision needed to connect these latent structural cues to quantitative property outcomes. Among the models, LLaMA-3 showed the most distinct clustering by crystal system even before finetuning, particularly in distinguishing high-symmetry groups like cubic and trigonal. RoBERTa and T5 also demonstrate some clustering by structural categories, albeit with greater overlap and more diffuse boundaries. These patterns indicate that pretrained language models are capable of learning latent structural relationships from text, despite not being trained specifically for materials science.
Once finetuned on the band gap prediction task, however, the clustering behavior undergoes a clear transition: embeddings now group primarily by band gap values rather than by crystal system, as visualized in Figure 4. This shift reveals that the fine-tuning process successfully reorients the model’s latent space to emphasize property-relevant dimensions over purely structural ones. The result is an embedding space where materials with similar band gaps are positioned closer together, regardless of their crystallographic classification. This transformation illustrates the effectiveness of supervised training in realigning general-purpose language representations toward targeted scientific objectives.
4 Conclusion
In this study, we investigated the use of transformer-based language models, RoBERTa, T5, and LLaMA-3, for predicting the band gaps of semiconductor materials directly from textual inputs. We demonstrated that these models can learn meaningful structure–property relationships without relying on handcrafted features or complex numerical descriptors.
Finetuned language models consistently outperformed shallow ML baselines, with LLaMA-3 achieving the highest accuracy using structured inputs (MAE 0.248 eV, $R^{2}$ 0.891). Even with natural language descriptions, the models captured relevant patterns, highlighting their flexibility for scenarios lacking structured data. Additionally, layer freezing analysis showed that strong predictive performance can be achieved with minimal finetuning. These findings underscore the adaptability of language models for scientific regression tasks.
Overall, our results establish transformer-based language models, particularly compact, decoder-only architectures like LLaMA-3, as promising tools for text-driven materials property prediction, offering scalability, interpretability, and efficiency for future materials informatics pipelines.
Code Availability Statement
[…]

### v3

@@
 Abstract
-We investigate the use of transformer-based language models, RoBERTa, T5, and LLaMA, for predicting the band gaps of semiconductor materials directly from textual representations that encode key material features such as chemical composition, crystal system, space group, number of atoms per unit cell, valence electron count, and other relevant electronic and structural properties.
-Quantum chemistry simulations such as Density Functional Theory (DFT) provide accurate predictions but are computationally intensive, limiting their feasibility for large-scale materials screening.
-Shallow machine learning (ML) models offer faster alternatives but typically require extensive data preprocessing to convert non-numerical material features into structured numerical inputs, often at the cost of losing critical descriptive information.
-In contrast, our approach leverages pretrained language models to process textual data directly, eliminating the need for manual feature engineering.
-We construct material descriptions in two formats: structured strings that combine key features in a consistent template, and natural language narratives generated using the ChatGPT API.
-For each model, we append a custom regression head and perform task-specific finetuning on a curated dataset of inorganic compounds.
-Our results show that finetuned language models, particularly the decoder-only LLaMA-3 architecture, can outperform conventional approaches in prediction accuracy and flexibility, achieving a mean absolute error (MAE) of 0.248 eV, root mean squared error (RMSE) of 0.345 eV, and $R^{2}$ of 0.891, compared to the best shallow ML baseline (XGBoost), which achieved an MAE of 0.318 eV, RMSE of 0.537 eV, and $R^{2}$ of 0.838.
-Notably, LLaMA-3 achieves competitive accuracy with minimal fine-tuning, suggesting its architecture enables more transferable representations for scientific tasks.
-This work demonstrates the effectiveness of finetuned language models for scientific property prediction and provides a scalable, language-native framework for materials informatics.
+We investigate transformer-based language models, including RoBERTa, T5, Llama-3, and MatSciBERT, for predicting the band gaps of semiconductor materials directly from textual descriptions.
+The inputs encode key material features, such as chemical composition, crystal system, space group, and other structural and electronic properties.
+Unlike shallow machine learning models, which require extensive feature engineering, or Graph Neural Networks, which rely on graph representations derived from atomic coordinates, pretrained language models can process textual inputs directly, eliminating the need for manual feature preprocessing or structure-based encoding.
+Material descriptions were constructed in two formats: structured strings with a consistent template and natural language narratives generated via the ChatGPT API.
+Each model was augmented with a custom regression head and finetuned for band gap prediction task.
+Language models of different architectures and parameter sizes were all able to predict band gaps from human-readable text with strong accuracy, achieving MAEs in the range of 0.25–0.33 eV, highlighting the success of this approach for scientific regression tasks.
+Finetuned Llama-3, with 1.2 billion parameters, achieved the highest accuracy (MAE 0.248 eV, $R^{2}$ 0.891).
+MatSciBERT, pretrained on materials science literature, reached comparable performance (MAE 0.288 eV, $R^{2}$ 0.871) with significantly fewer parameters (110 million), emphasizing the importance of domain-specific pretraining.
+Attention analysis shows that both models selectively focus on compositional and spin-related features while de-emphasizing geometric features, reflecting the difficulty of capturing spatial information from text.
+These results establish that pretrained language models can effectively extract complex feature-property relationships from textual material descriptions, providing a scalable, language-native framework for materials informatics.
 keywords:
 American Chemical Society,
+[…]
 […]
@@
 This reliance on extensive preprocessing and feature engineering not only adds complexity but also risks discarding nuanced or qualitative information, such as symmetry, bonding environments, or textual metadata, that could be valuable for accurate property prediction 10, 11.
-Recent advances in deep learning, particularly Graph Neural Networks (GNNs) 12, have improved the ability to capture the interconnectivity of atomic structures by representing them as graphs.
-However, this approach requires the conversion of atomic arrangements into graph representations, which adds another layer of preprocessing.
+Recent advances in deep learning, particularly Graph Neural Networks (GNNs) 12, have significantly enhanced the modeling of structure–property relationships by representing materials as atomic graphs that encode interatomic connectivity.
+However, these methods inherently rely on explicit structural information and require the conversion of atomic configurations into graph representations, introducing additional preprocessing steps and limiting their applicability to systems with well-defined crystal structures.
 Furthermore, GNNs still face limitations in integrating non-numerical properties, such as compound names, into the training process without additional preprocessing steps.
@@
 Additionally, CatBERTa, a RoBERTa-based predictive model, has been developed to predict adsorption energy in catalyst systems 13, 14.
-In this study, we explore the use of transformer-based language models, RoBERTa, T5, and LLaMA-3, to predict the band gaps of semiconductor materials directly from textual descriptions.
+In this study, we explore the use of transformer-based language models, RoBERTa, T5, Llama-3, and MatSciBERT to predict the band gaps of semiconductor materials directly from textual descriptions.
 These models enable the direct encoding of structured or natural language representations of materials, such as chemical composition, crystal symmetry, and electronic features, without requiring conventional feature engineering.
@@
 This approach allows models to learn mappings from text-based input to scalar band gap values.
-These models provide a flexible and generalizable framework for property prediction from textual materials data, extending the application of language models beyond conventional natural language tasks into scientific domains such as materials informatics.
+These models provide a flexible framework for property prediction from textual materials data, extending the application of language models beyond conventional natural language tasks into scientific domains such as materials informatics.
 2 Methods
 2.1 RoBERTa
-RoBERTa (Robustly Optimized BERT Approach) is a transformer-based encoder-only language model that builds upon the BERT architecture by refining pretraining strategies and removing restrictive design choices 20.
-It is widely recognized for its ability to capture rich contextual relationships between tokens through deep bidirectional self-attention, making it highly effective for a range of natural language processing tasks.
-RoBERTa is pretrained using the masked language modeling (MLM) objective, where a subset of tokens is masked and the model learns to predict them based on surrounding context.
-Unlike BERT, RoBERTa omits the next-sentence prediction task and instead trains on longer, contiguous sequences of text 20.
-It also introduces dynamic masking, larger batch sizes, and improved sentence packing strategies, which collectively enhance the generality and robustness of the learned representations.
-Tokenization is performed using a byte-level Byte-Pair Encoding (BPE) tokenizer, which allows the model to effectively handle multilingual text, rare subwords, and non-standard Unicode characters.
-In this study, we leverage the RoBERTa-base variant, comprising 12 transformer layers, 12 attention heads per layer, and a hidden dimension of 768.
-While RoBERTa provides strong pretrained representations, our framework emphasizes the necessity of objective-specific fine-tuning to adapt the model to domain-specific tasks such as band gap prediction.
-The model’s encoder-only structure offers advantages in processing structured text inputs efficiently, particularly in cases where outputs are scalar values rather than generated sequences.
+RoBERTa (Robustly Optimized BERT Approach) is an encoder-only transformer that refines BERT’s pretraining strategy through dynamic masking and longer sequence training 20.
+In this work, we employ the RoBERTa model (12 layers, 12 attention heads, 768 hidden units) as one of the backbone architectures for band gap prediction.
+Its bidirectional attention enables effective contextual encoding of structured text inputs derived from material features.
+Rather than emphasizing linguistic tasks, our framework fine-tunes RoBERTa to learn correlations between textual feature descriptions and scalar band gap values.
+This objective-specific adaptation allows the model to repurpose general language representations for scientific property prediction.
 2.2 T5
-T5 (Text-to-Text Transfer Transformer) is a unified encoder-decoder architecture that reframes all NLP tasks—including classification, translation, and regression—as text-to-text problems 21.
-In this study, we utilize the T5-small variant, which consists of 6 encoder and 6 decoder layers, each with a hidden size of 512 and 8 attention heads.
-Unlike encoder-only models such as RoBERTa, T5 processes input and output sequences through separate but coordinated transformer stacks, allowing for flexible sequence-to-sequence modeling and richer bidirectional context capture within the encoder.
-T5 is pretrained using a span corruption objective, in which contiguous spans of tokens in the input are replaced by unique sentinel tokens, and the model learns to reconstruct the missing content.
-This span-level objective fosters stronger global reasoning compared to token-level masking and supports robust performance on tasks requiring understanding of longer context windows.
-Architectural enhancements include the use of relative positional embeddings rather than absolute positions, parameter sharing between input and output embeddings, and simplified layer normalization without additive bias.
-These design choices collectively reduce model complexity and improve generalization.
-Tokenization in T5 is performed using SentencePiece, a subword tokenizer that supports language-agnostic processing and eliminates the need for whitespace-based token segmentation 22, 23.
-It is particularly useful for scientific or symbolic inputs, as it can encode rare tokens and domain-specific terms with high fidelity.
-During training, subword regularization introduces stochasticity in tokenization, enhancing robustness by exposing the model to a diverse range of token combinations.
-While T5’s encoder-decoder structure is typically leveraged for generative tasks, it is also adaptable for scalar prediction tasks like band gap regression.
-In our implementation, we repurpose the encoder to process material descriptions and append a custom regression head to the encoder output.
+T5 (Text-to-Text Transfer Transformer) is a unified encoder–decoder architecture that reformulates all NLP tasks as text-to-text transformations 21.
+We employ the T5-small (6 encoder and 6 decoder layers, 512 hidden size, 8 attention heads) to explore how a sequence-to-sequence framework can be applied to scientific regression problems.
+Throughout this manuscript, we refer to this model simply as T5.
+Pretrained with a span corruption objective, T5 learns to reconstruct masked text segments, promoting stronger global reasoning than token-level masking.
+In our framework, we utilize only the encoder output and append a custom regression head for band gap prediction.
+This setup leverages T5’s contextual encoding ability while avoiding unnecessary generative complexity, allowing efficient adaptation to structured material descriptions and scalar property prediction.
 2.3 Llama-3
-LLaMA-3.2-1B is a lightweight, decoder-only transformer model derived from Meta’s LLaMA-3 family of large language models 24, 25.
-Designed to provide a strong performance-efficiency trade-off, this version incorporates approximately 1 billion parameters and benefits from structured pruning and knowledge distillation from larger models.
-The architecture emphasizes both representational depth and inference efficiency, making it suitable for scalable applications beyond conventional text generation tasks.
-Unlike encoder-only models like RoBERTa or encoder-decoder frameworks such as T5, LLAMA-3 operates in an autoregressive, decoder-only configuration.
-Despite its unidirectional nature, LLAMA-3 can learn rich representations through deeply stacked attention layers and efficient architectural enhancements.
-The model employs several innovations that distinguish it from earlier transformers: it uses RMSNorm for pre-normalization to improve training stability, SwiGLU as a more expressive activation function, and Rotary Positional Embeddings (RoPE) to better generalize to longer sequences 26.
-In some variants, Grouped Query Attention (GQA) is used to reduce computational overhead while maintaining strong modeling capacity.
-LLaMA-3 tokenization is performed using BPE via SentencePiece 22, which enables robust handling of diverse character sets and subword units.
-Numerals are decomposed into individual digit tokens, and rare or unknown characters are represented using byte-level encodings.
-These choices ensure flexibility in modeling scientific texts that may include specialized symbols, units, or chemical formulas.
-Although LLAMA-3 was originally developed for autoregressive text generation, its modular architecture and robust token-level representation capabilities make it a versatile candidate for regression tasks.
-The model is particularly well-suited for structured and semi-structured textual input, where consistent formatting allows it to encode domain-relevant relationships.
-In our context, a custom regression head is appended to the decoder’s final hidden states to adapt the model for scalar band gap prediction.
-As with other language models in this study, fine-tuning on task-specific objectives is necessary to align the model’s pretrained linguistic knowledge with the specialized needs of materials property prediction.
+Llama-3 is a family of decoder-only transformer models developed by Meta, designed to provide state-of-the-art performance with efficient scaling across parameter sizes 22, 23.
+In this work, we specifically use the lightweight Llama-3.2-1B variant (approximately one billion parameters, embedding size 2048), which offers a strong balance between performance and computational efficiency.
+Throughout this manuscript, we refer to this model simply as Llama-3.
+We adapt Llama-3 by attaching a custom regression head to the decoder output for band gap prediction.
+Architectural components such as RMSNorm, SwiGLU activation, and Rotary Positional Embeddings improve training stability and contextual reasoning, while the SentencePiece-based tokenizer effectively encodes scientific symbols and numerical expressions.
+This configuration enables Llama-3 to extend its pretrained linguistic representations toward modeling structured material descriptions for quantitative property prediction.
+2.4 MatSciBERT
+MatSciBERT is a BERT-based language model pretrained on materials science literature to capture domain-specific terminology, chemical formulas, and structural descriptions 24.
+It uses the standard BERT-base architecture with 12 encoder layers, 768-dimensional embeddings, 12 attention heads, and a 3072-dimensional feed-forward layer, while its masked language modeling pretraining is optimized for the symbolic and long-tail vocabulary common in scientific texts.
+This provides stronger representations for materials-related language than general-purpose models.
+In our framework, we employ the MatSciBERT encoder with a custom regression head to predict band gaps from structured material descriptions.
+This design leverages MatSciBERT’s domain-aware embeddings while maintaining architectural simplicity, allowing efficient adaptation to property prediction tasks within materials informatics.
+2.5 Shallow ML Models
+We implemented three conventional regression algorithms, Random Forest (RF), Support Vector Regression (SVR), and Extreme Gradient Boosting (XGBoost), using the same dataset of structured material descriptors.
+All models were trained to predict the band gap values from numerical features extracted from the materials database.
+The RF model was implemented using scikit-learn’s RandomForestRegressor.
+The model constructs an ensemble of decision trees trained on randomly sampled subsets of both data and features, and outputs the average prediction across all trees.
+A five-fold cross-validated grid search was used to optimize key hyperparameters.
+The final configuration employed 1,000 trees, a maximum depth of 50, a minimum of two samples per split, and two samples per leaf.
+For the SVR, we used scikit-learn’s SVR implementation with a radial basis function (RBF) kernel.
+The model projects the input features into a high-dimensional space where a linear regression is performed within an $\epsilon$-insensitive margin.
+Grid search with five-fold cross-validation determined the optimal hyperparameters: a penalty parameter $C=5000$, $\epsilon=0.1$, and an RBF kernel.
+The high value of $C$ enables the model to fit complex nonlinear patterns, while $\epsilon$ controls the tolerance for prediction deviations near the regression boundary.
+The XGBoost model was implemented using the XGBRegressor from the xgboost library.
+XGBoost builds an ensemble of boosted decision trees, where each successive tree is trained to correct the residual errors of the previous ensemble.
+Grid search with five-fold cross-validation identified the best configuration: 2,000 trees, maximum depth of 9, learning rate of 0.1, subsample ratio of 1.0, and column sampling ratio of 0.6.
+These settings yield deep, expressive trees while maintaining regularization through feature subsampling.
+The model objective was set to minimize the squared error loss.
 2.6 Dataset
-In this study, we utilized the AFLOW database, a comprehensive open repository for computational materials science that contains extensive information on inorganic crystalline materials and their properties 27, 28.
+In this study, we utilized the AFLOW database, a comprehensive open repository for computational materials science that contains extensive information on inorganic crystalline materials and their properties 25, 26.
 Band gap calculations in AFLOW combine first-principles methods with empirical corrections through an automated workflow.
 The framework uses VASP to perform DFT calculations with the GGA-PBE functional for standard compounds while applying the GGA+U method for strongly correlated systems containing d- and f-block elements.
-To address GGA’s tendency to underestimate band gaps, AFLOW employs an empirical correction scheme based on a linear regression model derived from 102 benchmark compounds with known experimental values29, 27, 30.
+To address GGA’s tendency to underestimate band gaps, AFLOW employs an empirical correction scheme based on a linear regression model derived from 102 benchmark compounds with known experimental values27, 25, 28.
 This systematic approach, along with the database’s vast size and rich feature space, makes AFLOW particularly well-suited for ML tasks aimed at material property prediction.
@@
 The lower bound of 0 eV represents materials with metallic behavior, where there is no electronic band gap.
-Materials that have band gap higher than 5 eV are insulating materials, which have too large band gaps and do not conduct electricity under normal conditions 31, 3.
+Materials that have band gap higher than 5 eV are insulating materials, which have too large band gaps and do not conduct electricity under normal conditions 29, 3.
 By focusing on materials within this range of 0-5 eV band gaps, we ensure that our model targets materials with practical applications in electronics and optoelectronics.
@@
 Prompts were configured to ensure descriptions remained within a 512-token limit to maintain compatibility with the tokenizer constraints of RoBERTa, and to accommodate input length limits for T5 and Llama-3 as well.
-Both formats were applied uniformly across all three models: RoBERTa, T5, and LLaMA-3.
+Both formats were applied uniformly across all three models: RoBERTa, T5, Llama-3, and MatSciBERT.
 The models processed these inputs through their native tokenization pipelines, without additional handcrafted feature engineering.
-For each model, we appended a custom regression head to enable scalar prediction of band gap values and performed fine-tuning on the downstream task using the corresponding textual inputs.
 2.8 Input Features
@@
 The chemical formula represents the basic building blocks of the material, providing critical information about its stoichiometry and composition.
-The nature and type of atoms constituting the material greatly influence its electronic properties32, 9, 33.
-For example, the number of valence electrons of each species is crucial for band gap predictions34.
-The total number of valence electrons per unit cell helps predict how these electrons will contribute to the band structure, which is directly related to the band gap.
+The nature and type of atoms constituting the material greatly influence its electronic properties30, 9, 31.
+Specifically, the elemental types and their ratios determine atomic energy levels, such as s, p, and d orbitals, as well as bond types like covalent and ionic bonds, and electronegativity differences.
+These factors affect the relative positions and separation of the valence and conduction bands.
+In ionic compounds, a larger electronegativity difference leads to a greater energy separation between the valence band, which is primarily formed by anion orbitals, and the conduction band, which is mainly formed by cation orbitals.
+This typically results in a larger band gap.
+For example, II-VI compounds often have larger band gaps compared to III-V compounds32.
+For example, the number of valence electrons of each species is crucial for band gap predictions33.
+The total number of valence electrons per unit cell governs how the electronic bands are filled and directly influences the position of the Fermi level34.
+Differences in valence electron count and orbital configuration among elements or compounds can lead to significant variations in both the magnitude and nature of the band gap.
+Materials with stronger orbital overlap and higher structural symmetry tend to have wider bands and smaller, often direct, band gaps, whereas weaker overlap or lower symmetry can result in indirect or larger gaps, ultimately determining the optical and transport properties of the material35.
 Additionally, we paid particular attention to structural features, including crystal class, family, and system, as well as lattice parameters, which include the dimensions and angles of the unit cell.
-These factors not only shape the arrangement of atoms and their interactions but also define the symmetry and geometric properties of the crystal, directly influencing the distribution of electronic states within the energy bands 35, 36, 30.
+These factors not only shape the arrangement of atoms and their interactions but also define the symmetry and geometric properties of the crystal, directly influencing the distribution of electronic states within the energy bands 36, 37, 28.
+The crystal’s symmetry determines the shape of the Brillouin zone and band degeneracy.
+High symmetry, as seen in cubic systems, often leads to high band degeneracy at high-symmetry points, where multiple electrons have the same energy, resulting in a relatively simple band structure.
+In contrast, lower symmetry, as in orthorhombic systems, can cause band splitting, which may change the band gap’s size and type38.
 The space group and point group information were included to account for the effects of symmetry on electronic states and band splitting.
@@
 c Visualization of the finetuning process.
-The input text undergoes tokenization and embedding through a pretrained LLM, followed by a custom regression head for prediction.
+The input text undergoes tokenization and embedding through multiple model architectures (RoBERTa, T5, Llama-3, MatSciBERT), followed by a custom regression head for prediction.
 d Demonstrates the Transformer encoder and the multi-head attention mechanism with Query (Q), Key (K), and Value (V) operations.
-We developed a language model-based framework to predict the band gaps of semiconductor materials using transformer-based language models: RoBERTa, T5, and LLaMA-3.
+We developed a language model-based framework to predict the band gaps of semiconductor materials using transformer-based language models: RoBERTa, T5, Llama-3, and MatSciBERT.
 As shown in Figure 1a, we constructed the dataset from the AFLOW database by extracting relevant material features and transforming them into textual formats suitable for language model inputs.
@@
 These formats enabled us to assess how each model handles both highly regular and semantically rich input styles.
-Additional examples of both the structured strings and natural language descriptions are provided in the Supporting Information.s
-At the core of our approach are transformer-based language models, each differing in architecture and tokenization strategy.
-While RoBERTa uses an encoder-only structure, T5 adopts an encoder-decoder design, and LLaMA-3 operates with a decoder-only architecture.
+Additional examples of both the structured strings and natural language descriptions are provided in the Supporting Information.
+At the core of our approach are transformer-based language models, each differing in architecture, tokenization strategy, and training strategy.
+Our approach is built on transformer-based language models with distinct architectures and training schemes.
+RoBERTa and MatSciBERT share an encoder-only design, but MatSciBERT is additionally pretrained on materials science texts for domain-specific understanding.
+T5 uses an encoder–decoder structure, while Llama-3 is a decoder-only model.
 Each model tokenizes the input text using its native tokenizer, such as byte-level BPE or SentencePiece, and processes the sequence to generate contextual embeddings (Figure 1c).
-These embeddings are passed through a custom regression head to produce a scalar band gap prediction, as shown in Figure 1d.
-The design of this framework allows for direct comparison across model types and input formats while minimizing the need for handcrafted features or domain-specific encodings.
+These embeddings are passed through a custom regression head to produce a scalar band gap prediction.
 3.2 Model Performance
-We evaluated the performance of three transformer-based language models, RoBERTa, T5, and LLaMA-3, on the task of predicting semiconductor band gaps from text-based material descriptions.
-Each model was finetuned with a custom regression head and assessed using two types of input formats: a structured string representation and a natural language description generated via GPT-3.5 Turbo.
-These models were compared against shallow ML baselines, which also use the same input data format), including support vector regression (SVR), random forest, and XGBoost.
-The accuracy of the model was quantified using three metrics: The mean absolute error (MAE), the root mean square error (RMSE), and the coefficient of determination ($R^{2}$), as shown in Table 2, with the parity plots visualized in Figure 2.
+We evaluated the performance of four transformer-based language models—RoBERTa, T5, Llama-3, and MatSciBERT—on the task of predicting semiconductor band gaps from text-based material descriptions.
+We also included shallow machine learning baselines using the same input format, including SVR, random forest, and XGBoost.
+GNNs were not considered, as their graph-based input representations do not align with the feature-based approach used here.
+Model accuracy was quantified using three metrics: mean absolute error (MAE), root mean square error (RMSE), and the coefficient of determination ($R^{2}$).
+The results are summarized in Table 2, with parity plots shown in Figure 2.
 Table 2: Comparison of model performance across different ML approaches.
-Each transformer-based model was evaluated using both structured string and description input formats.
-Best performance per metric is shown in bold.
+For LLMs, results are shown as structured string / natural language description.
+Best performance per metric is shown in bold, and second-best performance is underlined.
 Model Model Type MAE (eV) RMSE (eV) $R^{2}$
@@
 Llama-3(description) LLM (Decoder) $0.335\pm 0.008$ $0.473\pm 0.006$ $0.843\pm 0.054$
-Figure 2: Parity plots for band gap predictions across models: a SVR, b XGBoost, c Random Forest, d RoBERTa, e T5, f LLaMA-3
-Across all models, inputs in structured string format consistently outperformed descriptive natural language inputs.
-Among the transformer-based models, LLaMA-3 with structured input achieved the best overall performance, with a mean absolute error of 0.248 eV and R² of 0.891.
+MatSciBERT(string) LLM (Encoder) $0.288\pm 0.007$ $0.407\pm 0.006$ $0.871\pm 0.004$
+MatSciBERT(description) LLM (Encoder) $0.366\pm 0.009$ $0.503\pm 0.007$ $0.808\pm 0.007$
+Figure 2: Parity plots for band gap predictions across models: a SVR, b XGBoost, c Random Forest, d RoBERTa, e T5, f Llama-3, g MatSciBERT
+Across all models, inputs in structured string format consistently outperformed descriptive natural language inputs, as shown in Table 2.
+Both input types are derived from the same underlying features, so the richness of information is equivalent; the difference arises from how the information is presented.
+Structured strings provide a consistent, systematic representation that reduces variability, making it easier for the model to learn direct mappings between features and target properties.
+In contrast, language descriptions introduce variability in phrasing, terminology, and length, requiring the model to first infer the relevant features from context.
+This adds complexity and can obscure patterns, even though the same information is present.
+Among the transformer-based models, Llama-3 with structured input achieved the best overall performance, with a mean absolute error of 0.248 eV and $R^{2}$ of 0.891 (Table 2).
+Following closely, MatSciBERT delivered a strong performance with an MAE of 0.288 eV and an $R^{2}$ of 0.872.
 While RoBERTa and T5 also outperformed the shallow ML models, their performance was lower when using GPT-generated descriptions compared to structured string inputs.
-However, the ability of all three language models to still learn meaningful correlations from unstructured natural language input underscores a key strength of this approach.
-Rather than viewing this variability as a limitation, we interpret it as evidence of the models’ capacity to generalize from flexible, human-like descriptions of materials.
-Notably, even in its smallest size, LLaMA-3 demonstrated strong capacity for property prediction when paired with fine-tuning, indicating that decoder-only architectures can serve as lightweight yet powerful alternatives for scientific regression tasks.
+These results highlight several insights.
+First, language models can predict band gaps directly from text-based inputs without any feature preprocessing.
+Given its large number of parameters, it is not surprising that Llama-3 performed the best.
+More interestingly, when comparing RoBERTa and MatSciBERT, which share similar architectures, MatSciBERT achieved better results despite having fewer parameters.
+This suggests that the training strategy and the use of a domain-specific corpus are critical.
+Pretraining on materials science literature enables the model to more effectively capture the relationship between material features and band gap values.
+Even though we evaluated multiple types of language models, it is difficult to generalize performance purely based on architecture since the number of parameters varies greatly.
+Nonetheless, it is important to demonstrate that text-based band gap prediction works across different model types.
+It is also not surprising that language model approaches outperform shallow machine learning methods.
+To clarify, both shallow ML models and language models use the same feature set, so the richness of information is identical.
+The superior performance of language models indicates their enhanced capability to learn complex correlations between the provided features and the band gap.
 3.3 Layer Freezing Analysis
-To assess the extent to which each model’s pretrained representations contribute to band gap prediction, we conducted layer freezing experiments for RoBERTa, T5, and LLAMA-3 using structured string inputs.
-In each experiment, we progressively froze layers and finetuned only a portion of the model.
-The goal was to determine how much domain-specific tuning is necessary to achieve high predictive accuracy.
-Table 3 report the MAE, RMSE, and $R^{2}$ values for each model under various freezing configurations.
-In all cases, a custom regression head was trained on top of the partially or fully frozen transformer backbone.
-Table 3: Comparison of layer freezing strategies across RoBERTa, T5, and LLaMA-3 using structured string inputs.
+To investigate how pretrained representations in each model contribute to band gap prediction, we conducted layer-freezing experiments on RoBERTa, T5, Llama-3, and MatSciBERT using structured string inputs.
+In these experiments, we progressively froze layers of the transformer backbone and fine-tuned only a portion of the model, while training a custom regression head on top of the frozen or partially frozen representations.
+The goal was not to reduce computational cost, but to probe which layers carry the most task-relevant information.
+Table 3 reports the MAE, RMSE, and $R^{2}$ values for each model under various freezing setups.
+This analysis provides insight into how much domain-specific fine-tuning is necessary and which parts of the pretrained model are most informative for the downstream task, highlighting the layers that contribute most strongly to predictive accuracy.
+Table 3: Comparison of layer freezing strategies across RoBERTa, T5, Llama-3, and MatSciBERT using structured string inputs.
 The fully finetuned (non-frozen) results are included for reference.
+Parameter counts and percentages represent the number of trainable parameters relative to the no-freezing case.
 All values are reported as mean $\pm$ standard deviation.
-Model Freezing Strategy MAE (eV) RMSE (eV) $R^{2}$
-RoBERTa Fully finetuned (no freezing) $0.325\pm 0.006$ $0.447\pm 0.005$ $0.855\pm 0.004$
-Freeze first layer $0.328\pm 0.008$ $0.448\pm 0.006$ $0.848\pm 0.005$
-Freeze all but final 3 layers $0.388\pm 0.009$ $0.510\pm 0.007$ $0.817\pm 0.006$
-Freeze all but final layer $0.509\pm 0.012$ $0.648\pm 0.009$ $0.721\pm 0.009$
-Freeze all layers $0.663\pm 0.016$ $0.826\pm 0.011$ $0.563\pm 0.013$
-T5 Fully finetuned (no freezing) $0.301\pm 0.007$ $0.448\pm 0.006$ $0.861\pm 0.005$
-Freeze first layer $0.350\pm 0.008$ $0.504\pm 0.007$ $0.849\pm 0.006$
-Freeze all but final 3 layers $0.367\pm 0.009$ $0.516\pm 0.007$ $0.832\pm 0.006$
-Freeze all but final layer $0.598\pm 0.014$ $0.784\pm 0.011$ $0.619\pm 0.011$
-Freeze all layers $0.792\pm 0.019$ $0.981\pm 0.013$ $0.420\pm 0.014$
-LLaMA-3 Fully finetuned (no freezing) $\mathbf{0.248\pm 0.006}$ $\mathbf{0.345\pm 0.005}$ $\mathbf{0.891\pm 0.004}$
-Freeze first layer $0.279\pm 0.007$ $0.426\pm 0.006$ $0.878\pm 0.004$
-Freeze all but final 3 layers $0.318\pm 0.008$ $0.474\pm 0.006$ $0.851\pm 0.005$
-Freeze all but final layer $0.424\pm 0.010$ $0.576\pm 0.008$ $0.793\pm 0.007$
-Freeze all layers $0.716\pm 0.017$ $0.893\pm 0.012$ $0.518\pm 0.013$
+Model Freezing Strategy Parameters (count, %) MAE (eV) RMSE (eV) $R^{2}$
+RoBERTa Fully finetuned (no freezing) 125,236,993 (100%) $0.325\pm 0.006$ $0.447\pm 0.005$ $0.855\pm 0.004$
+Freeze first layer 118,149,121 (94.3%) $0.328\pm 0.008$ $0.448\pm 0.006$ $0.848\pm 0.005$
+Freeze all but final 3 layers 61,446,145 (49.1%) $0.388\pm 0.009$ $0.510\pm 0.007$ $0.817\pm 0.006$
+Freeze all but final layer 8,269,825 (6.6%) $0.509\pm 0.012$ $0.648\pm 0.009$ $0.721\pm 0.009$
+Freeze all layers 591,361 (0.5%) $0.663\pm 0.016$ $0.826\pm 0.011$ $0.563\pm 0.013$
+T5 Fully finetuned (no freezing) 60,769,793 (100%) $0.301\pm 0.007$ $0.448\pm 0.006$ $0.861\pm 0.005$
+Freeze first layer 57,622,785 (94.8%) $0.350\pm 0.008$ $0.504\pm 0.007$ $0.849\pm 0.006$
+Freeze all but final 3 layers 21,830,401 (19.8%) $0.367\pm 0.009$ $0.516\pm 0.007$ $0.832\pm 0.006$
+Freeze all but final layer 7,671,553 (7.0%) $0.598\pm 0.014$ $0.784\pm 0.011$ $0.619\pm 0.011$
+Freeze all layers 592,129 (0.5%) $0.792\pm 0.019$ $0.981\pm 0.013$ $0.420\pm 0.014$
+Llama-3 Fully finetuned (no freezing) 1,237,915,649 (100%) $\mathbf{0.248\pm 0.006}$ $\mathbf{0.345\pm 0.005}$ $\mathbf{0.891\pm 0.004}$
+Freeze first layer 1,177,094,145 (95.1%) $0.279\pm 0.007$ $0.426\pm 0.006$ $0.878\pm 0.004$
+Freeze all but final 3 layers 447,236,097 (36.1%) $0.318\pm 0.008$ $0.474\pm 0.006$ $0.851\pm 0.005$
+Freeze all but final layer 325,593,089 (26.3%) $0.424\pm 0.010$ $0.576\pm 0.008$ $0.793\pm 0.007$
+Freeze all layers 2,101,249 (0.2%) $0.716\pm 0.017$ $0.893\pm 0.012$ $0.518\pm 0.013$
+MatSciBERT Fully finetuned (no freezing) 110,509,825 (100%) $0.288\pm 0.007$ $0.407\pm 0.006$ $0.871\pm 0.004$
+Freeze first layer 103,421,953 (93.6%) $0.293\pm 0.007$ $0.404\pm 0.005$ $0.874\pm 0.004$
+Freeze all but final 3 layers 22,445,569 (20.3%) $0.340\pm 0.008$ $0.454\pm 0.006$ $0.849\pm 0.005$
+Freeze all but final layer 8,269,825 (7.5%) $0.415\pm 0.010$ $0.539\pm 0.007$ $0.798\pm 0.007$
+Freeze all layers 1,181,953 (1.1%) $0.827\pm 0.020$ $0.994\pm 0.013$ $0.385\pm 0.015$
 Throughout this study, we refer to the original language models, trained on general text corpora prior to any materials-specific adaptation, as “pretrained” models.
 After supervised training on the band gap prediction task, we refer to the models as “finetuned.”
-In all three models, predictive performance improved progressively as more transformer layers were unfrozen, highlighting the importance of task-specific fine-tuning in scientific regression applications.
-This trend highlights a key insight: while LLMs capture broadly useful representations during pretraining, these representations alone are insufficient to achieve optimal accuracy in domain-specific prediction tasks without additional adaptation.
-In the case of RoBERTa and T5, the most notable performance gains occurred when the final three layers of the model were unfrozen.
-This suggests that the highest-level contextual embeddings learned during pretraining are not directly aligned with the band gap target property but can be effectively adapted through partial retraining.
-LLaMA-3, on the other hand, exhibited relatively strong performance even when the majority of its layers were frozen.
-For example, when only the final three layers were unfrozen during fine-tuning, LLaMA-3 achieved $R^{2}$ values comparable to those of fully fine-tuned RoBERTa and T5 models.
-This observation points to its efficient architecture and inductive biases, which may make its learned representations more transferable to new scientific tasks.
-In particular, the decoder-only architecture of LLAMA-3, combined with its use of Rotary Positional Embeddings and SwiGLU activation functions, appears to confer greater flexibility in adapting to structured input formats such as the material strings used in this study.
-Furthermore, the ability of LLAMA-3 to achieve high $R^{2}$ values with only the final layer trainable supports its utility in resource-constrained scenarios, such as edge inference or few-shot scientific learning contexts.
-Collectively, these results reinforce the conclusion that while pretrained LLMs provide a valuable foundation, their effective use in materials science and other physical domains requires carefully calibrated fine-tuning strategies.
-Strategic layer freezing not only reduces computational burden during training but also offers insight into which parts of the model architecture contribute most directly to scientific prediction tasks.
-This understanding is critical for efficiently deploying transformer-based models in practical, data-limited, or performance-critical settings.
-3.4 Embedding Analysis
+In all four models, predictive performance improved progressively as more transformer layers were unfrozen, highlighting the importance of task-specific fine-tuning in scientific regression tasks, as shown in Table 3.
+While LLMs capture broadly useful representations during pretraining, these representations alone are insufficient to achieve optimal accuracy in domain-specific prediction tasks without additional adaptation.
+As more layers are involved in fine-tuning, the models can better capture feature–property correlations.
+Interestingly, MatSciBERT exhibits slightly better RMSE and $R^{2}$ scores compared to its fully finetuned case, which may be attributed to the well-formed representations learned from prior exposure to the materials science domain.
+Figure 3: Scaling behavior of finetuning strategies across transformer-based models.
+MAE is shown as a function of the number of trainable parameters.
+Colors indicate different model architectures: RoBERTa (yellow), T5 (green), Llama-3 (blue), and MatSciBERT (red).
+Marker shapes represent different freezing strategies, from fully finetuned (no freezing) to full layer freezing.
+The rough scaling relationship between the number of parameters and prediction accuracy is illustrated in Figure 3.
+Generally, as the number of trainable parameters increases, MAE decreases, indicating improved predictive performance across models and layer-freezing strategies.
+Within each model, this relationship is relatively linear in the layer-wise freezing experiments, showing that gradually unfreezing layers consistently improves performance.
+However, when comparing across different models, the relationship is not strictly linear.
+This reflects the fact that predictive performance depends not only on the number of parameters but also on other factors, including model architecture, pretraining corpus, task-specific fine-tuning strategies, and the alignment between learned representations and the structured input features.
+These results emphasize that while pretrained LLMs provide a strong foundation, achieving optimal performance in materials science requires careful selection of trainable layers and fine-tuning strategies, especially under computational or resource constraints.
+3.4 Feature-wise Self-Attention Score
+To identify which material features are most emphasized during band gap prediction, we conducted a feature-wise self-attention analysis on Llama-3 and MatSciBERT, the two best-performing models.
+In transformer architectures, the attention mechanism quantifies how strongly each token attends to others when forming contextual representations, as shown in Figure 1d.
+For each model, input text was tokenized (maximum length of 512 tokens, truncation enabled), and attention weights were extracted from the first and last transformer layers.
+The attention score here refers to the attention weight between the first token, used as the regression header, and all other tokens in the sequence.
+For each head, the attention scores from the regression token to the rest of the tokens were averaged, and the resulting values were then averaged across all heads within the layer.
+Attention scores between the regression header token and all feature tokens were extracted from each transformer layer.
+For a given layer $l$ and sample $i$, the attention weights were first averaged across all attention heads:
+$s_{i,l}(t)=\frac{1}{H}\sum_{h=1}^{H}s_{i,l,h}(t)$ (1)
+where $H$ is the number of attention heads and $s_{i,l,h}(t)$ denotes the attention score from the header token to token $t$ in head $h$.
+For each feature $f$, we identified its corresponding token span and recorded the maximum attention score within that span:
+$\text{score}_{i,l,f}^{\text{raw}}=\max_{t\in[t_{\text{start}},t_{\text{end}}]}s_{i,l}(t)$ (2)
+To ensure comparability across samples, raw scores were normalized to the range [0, 1] using min–max scaling:
+$\text{score}_{i,l,f}^{\text{norm}}=\frac{\text{score}_{i,l,f}^{\text{raw}}-\min_{k}\text{score}_{i,l,k}^{\text{raw}}}{\max_{k}\text{score}_{i,l,k}^{\text{raw}}-\min_{k}\text{score}_{i,l,k}^{\text{raw}}}$ (3)
+where $k$ indexes all features in sample $i$.
+Finally, layer-wise feature attention was obtained by averaging normalized scores across all test samples:
+$\text{Avg\_Attention}_{l,f}=\frac{1}{N}\sum_{i=1}^{N}\text{score}_{i,l,f}^{\text{norm}}$ (4)
+This procedure provides a consistent measure of how each layer attends to specific material features based on attention scores.
+It is important to note that these scores do not incorporate the value matrix; therefore, as illustrated in Figure 1d, they do not directly quantify feature importance.
+Instead, they offer a descriptive view of the model’s focus on the input tokens.
+Figure 4: Feature-wise self-attention scores for LLaMA-3 and MatSciBERT. a LLaMA-3, first layer; b LLaMA-3, final layer (layer 16); c MatSciBERT, first layer; d MatSciBERT, final layer (layer 12).
+The results reveal clear differences in attention patterns between the first and last layers for both models.
+In our analysis, the embedding of the first token from the last layer is ultimately used for regression, reflecting how the model allocates attention to optimize prediction performance.
+For LLaMA-3, attention to geometric features, such as lattice variation, lattice system, and overall geometry, decreases in the final layer, as seen by comparing Figure 4a and b.
+Although these features are important for structure–property relationships, this pattern suggests that LLaMA-3 struggles to directly extract geometric information from text-based descriptions, instead emphasizing compositional and spin-related features.
+MatSciBERT, which benefits from domain-specific pretraining on materials science literature, shows earlier specialization in crystallographic features, as illustrated in Figure 4c.
+Similar to LLaMA-3, attention to geometric features declines in later layers, highlighting the difficulty of encoding geometric information purely from text (Figure 4d).
+However, MatSciBERT distributes attention more broadly in the first layer, with structural symmetry attributes such as point group receiving relatively high scores.
+By the final layer, attention becomes more selective, focusing on key features such as composition, including species and compound information.
+3.5 Embedding Map
 We conducted embedding space analysis using t-SNE visualizations to investigate how each model organizes material representations before and after fine-tuning.
-As shown in Figure 3, we extracted the first-token embeddings from the pretrained and finetuned versions of RoBERTa, T5, and LLaMA-3 using structured string inputs, and colored the points by crystal system.
+As shown in Figure 5 and 6, we extracted the first-token embeddings from the pretrained and finetuned versions of RoBERTa, T5, Llama-3, and MatSciBERT using structured string inputs, and colored the points by crystal system.
 In the “pretrained” condition, which corresponds to the original language model weights before finetuning on the band gap prediction task, the embeddings reflect structural signals learned from general language corpora.
-Figure 3: t-sne visualizations of embedding analysis with crystal systems. a pretrained RoBERTa, b finetuned RoBERTa, c pretrained T5, d finetuned T5, e pretrained LLAMA-3, f finetuned LLaMa-3
-Figure 4: t-SNE visualizations of embedding analysis for bandgap. a pretrained RoBERTa, b finetuned RoBERTa, c pretrained T5, d finetuned T5, e pretrained LLaMa-3, f finetuned LLaMa-3
-In the pretrained state, the embeddings of all three models exhibit meaningful clustering based on the crystal system.
-This structure-aware behavior is likely driven by explicit features in the input, such as crystal class, space group, and lattice parameters, information that correlates with symmetry.
-However, while these features can separate materials by structural family, they do not offer sufficient signal to accurately predict electronic properties like band gap.
-Without finetuning, the models lack the task-specific supervision needed to connect these latent structural cues to quantitative property outcomes.
-Among the models, LLaMA-3 showed the most distinct clustering by crystal system even before finetuning, particularly in distinguishing high-symmetry groups like cubic and trigonal.
-RoBERTa and T5 also demonstrate some clustering by structural categories, albeit with greater overlap and more diffuse boundaries.
-These patterns indicate that pretrained language models are capable of learning latent structural relationships from text, despite not being trained specifically for materials science.
-Once finetuned on the band gap prediction task, however, the clustering behavior undergoes a clear transition: embeddings now group primarily by band gap values rather than by crystal system, as visualized in Figure 4.
-This shift reveals that the fine-tuning process successfully reorients the model’s latent space to emphasize property-relevant dimensions over purely structural ones.
-The result is an embedding space where materials with similar band gaps are positioned closer together, regardless of their crystallographic classification.
-This transformation illustrates the effectiveness of supervised training in realigning general-purpose language representations toward targeted scientific objectives.
+The “fine-tuned” condition represents embeddings generated after the models were fully fine-tuned for band gap prediction.
+Figure 5: t-SNE visualizations of embeddings colored by crystal system.
+a-d show results from the pretrained models: a RoBERTa, b T5, c Llama-3, d MatSciBERT.
+e-h show results from the corresponding finetuned models: e RoBERTa, f T5, g Llama-3, h MatSciBERT.
+Figure 6: t-SNE visualizations of embeddings for bandgap prediction.
+a-d show results from the pretrained models: a RoBERTa, b T5, c Llama-3, d MatSciBERT.
+e-h show results from the corresponding finetuned models: e RoBERTa, f T5, g Llama-3, h MatSciBERT.
+In the pretrained state, embeddings from all four models exhibit meaningful clustering according to crystal system, as shown in Figure 5a-d.
+In contrast, when coloring by bandgap (Figure 6a-d), no clear clustering is observed.
+This structure-aware behavior likely arises from explicit input features such as crystal class, space group, and lattice parameters, which correlate with symmetry.
+While these features can separate materials by structural family, they do not provide sufficient signal to predict electronic properties like bandgap.
+In other words, without finetuning, the models lack task-specific supervision to map structural cues to quantitative property outcomes.
+After finetuning on the bandgap prediction task, however, the clustering behavior changes markedly: embeddings now group primarily by bandgap rather than by crystal system, as visualized in Figure 5e-h and Figure 6e-h.
+In particular, Llama-3 shows a significant loss of crystal system clustering while forming a near-linear embedding map based on bandgap, which likely contributes to its superior predictive performance.
+The result is an embedding space where materials with similar bandgaps are positioned closer together, independent of their crystallographic classification.
+This transformation illustrates how supervised training can reorient general-purpose language representations toward property-specific scientific objectives.
+T5 produces a qualitatively different t-SNE embedding: unlike the roughly circular maps seen in most other models, T5’s embedding forms a curvy, elongated shape.
+We note that the models have differing embedding dimensions, and t-SNE can distort distances depending on dimensionality and distribution.
+Therefore, some apparent differences may reflect visualization artifacts or latent space geometry rather than intrinsic model behavior.
+Further quantitative analysis of cluster separation and latent structure would be needed to confirm these patterns.
 4 Conclusion
-In this study, we investigated the use of transformer-based language models, RoBERTa, T5, and LLaMA-3, for predicting the band gaps of semiconductor materials directly from textual inputs.
-We demonstrated that these models can learn meaningful structure–property relationships without relying on handcrafted features or complex numerical descriptors.
-Finetuned language models consistently outperformed shallow ML baselines, with LLaMA-3 achieving the highest accuracy using structured inputs (MAE 0.248 eV, $R^{2}$ 0.891).
+In this study, we investigated the use of transformer-based language models, RoBERTa, T5, Llama-3, and MatSciBERT for predicting the band gaps of semiconductor materials directly from textual inputs.
+We demonstrated that these models can learn meaningful feature–property relationships without relying on complex feature preprocessing or structure graph of atomic coordinates.
+Llama-3 achieving the highest accuracy using structured inputs (MAE 0.248 eV, $R^{2}$ 0.891).
 Even with natural language descriptions, the models captured relevant patterns, highlighting their flexibility for scenarios lacking structured data.
-Additionally, layer freezing analysis showed that strong predictive performance can be achieved with minimal finetuning.
-These findings underscore the adaptability of language models for scientific regression tasks.
-Overall, our results establish transformer-based language models, particularly compact, decoder-only architectures like LLaMA-3, as promising tools for text-driven materials property prediction, offering scalability, interpretability, and efficiency for future materials informatics pipelines.
+Overall, this work highlights the potential of both general-purpose and domain-specific language models as flexible, scalable, and efficient tools for materials property prediction.
+By enabling direct prediction of properties from human-readable text, these models remove the need for extensive feature engineering or graph-based structure encoding, allowing rapid, end-to-end property estimation from structured or natural language material descriptions.
+The results also demonstrate that pretrained language models can provide interpretable insights through attention and embedding analyses, identifying the material features most relevant to the target property.
+.
 Code Availability Statement
 […]
+Technology Use Disclosure
+We used ChatGPT and Claude to help with grammar and typographical corrections during the preparation of this manuscript.
+The authors have carefully reviewed, verified, and approved all content to ensure accuracy and integrity.